### v1

# Differential Privacy for Deep Learning in Medicine

Abstract
Differential privacy (DP) is a key technique for protecting sensitive patient data in medical deep learning (DL). As clinical models grow more data-dependent, balancing privacy with utility and fairness has become a critical challenge. This scoping review synthesizes recent developments in applying DP to medical DL, with a particular focus on DP-SGD and alternative mechanisms across centralized and federated settings. Using a structured search strategy, we identified 74 studies published up to March 2025. Our analysis spans diverse data modalities, training setups, and downstream tasks, and highlights the tradeoffs between privacy guarantees, model accuracy, and subgroup fairness. We find that while DP—especially at strong privacy budgets—can preserve performance in well-structured imaging tasks, severe degradation often occurs under strict privacy, particularly in underrepresented or complex modalities. Furthermore, privacy-induced performance gaps disproportionately affect demographic subgroups, with fairness impacts varying by data type and task. A small subset of studies explicitly addresses these tradeoffs through subgroup analysis or fairness metrics, but most omit them entirely. Beyond DP-SGD, emerging approaches leverage alternative mechanisms, generative models, and hybrid federated designs, though reporting remains inconsistent. We conclude by outlining key gaps in fairness auditing, standardization, and evaluation protocols, offering guidance for future work toward equitable and clinically robust privacy-preserving DL systems in medicine.
1 Introduction
Deep learning (DL) has become a cornerstone of modern medical artificial intelligence (AI), driving advances in diagnostic imaging, survival prediction, biosignal analysis, and electronic health record (EHR) modeling [1, 2, 3]. However, the deployment of DL models in clinical settings raises critical concerns regarding patient privacy [4]. Medical data are inherently sensitive, and DL models trained on such data are vulnerable to various privacy attacks—including membership inference, model inversion, and identity disclosure—potentially exposing the presence or characteristics of individual patients in the training set [5, 6, 7, 8, 9].
Differential privacy (DP) [10] provides formal guarantees to protect individuals within a dataset by bounding the influence of any single record on the model’s output. In DL, this is commonly implemented via differentially private stochastic gradient descent (DP-SGD) [11], which injects noise into clipped gradients during training (Fig. 1). Despite its theoretical strength, practical deployment of DP in healthcare DL introduces key trade-offs: noise reduces model utility and may disproportionately affect underrepresented subgroups, potentially reinforcing existing healthcare inequities [12, 13, 5].
To better understand the state of DP in clinical DL, we conduct a comprehensive scoping review of empirical studies that apply DP mechanisms in medical contexts. While several prior reviews address DP broadly or focus on applications in generic domains, no existing synthesis systematically evaluates how DP interacts with clinical data, model design, privacy-utility trade-offs, fairness impacts, or adversarial robustness.
To guide our review, we investigate the following research questions:
  • RQ1:
    To what extent do medical DL models leak sensitive information, and under what conditions are privacy breaches most likely?
  • RQ2:
    How effectively do DP-based methods mitigate membership inference and other privacy attacks in healthcare applications?
  • RQ3:
    What are the empirical impacts of DP on model performance, fairness, and clinical applicability in real-world scenarios?
This scoping review aims to critically synthesize current knowledge, identify key challenges and opportunities, and provide practical guidance for future research and clinical practice involving privacy-preserving DL models in medicine. In the following sections, we first review the applications of DP across clinical DL tasks, covering data modalities, training settings, and privacy accounting methods. We then analyze how architectural choices, normalization strategies, and pretraining affect the privacy-utility tradeoff, with a focus on model performance across varying privacy budgets. Next, we examine fairness impacts, highlighting subgroup disparities, the metrics used to evaluate them, and gaps in current practices. We also survey alternative privacy mechanisms beyond DP-SGD, including federated learning (FL), local perturbation, and hybrid methods. Finally, we synthesize empirical findings on privacy attacks and defenses, before concluding with a discussion of key limitations, open challenges, and recommendations for building robust, equitable, and privacy-preserving AI systems in medicine.
Figure 1: Overview of privacy risks and defenses in different phases of medical deep learning. (a) During model training, raw patient data—collected from clinical sources—can be vulnerable to leakage if accessed by unauthorized parties. Differential privacy is applied at this stage, typically by injecting noise into model gradients, to prevent individual data from being memorized. (b) After deployment, the trained model is exposed to potential inference attacks, such as membership inference, where adversaries aim to determine if specific individuals were in the training set. Models trained with differential privacy are more resistant to such attacks, enhancing patient confidentiality throughout the model lifecycle.
2 Methods
This scoping review was conducted in accordance with the Preferred Reporting Items for Systematic Reviews and Meta-Analyses extension for Scoping Reviews (PRISMA-ScR) guidelines [14, 15] to ensure methodological rigor and transparency [16]. The complete PRISMA-ScR checklist is provided in Supplementary Table 1. The methodological approach to data collection, article selection, and results synthesis is described in detail below.
2.1 Eligibility Criteria
Any study was eligible if it was published on or before March 1, 2025, and explicitly focused on empirical applications of DP in conjunction with DL methods for healthcare-related tasks. Eligible articles included original peer-reviewed journal papers, conference proceedings, and impactful preprints in English. Studies purely theoretical in nature or without explicit medical or clinical applications were excluded. Review articles, editorials, and commentaries were also excluded to focus exclusively on primary empirical contributions.
2.2 Information Sources
We systematically searched four electronic databases: PubMed, IEEE Xplore, ACM Digital Library, and Web of Science. Database queries were performed on March 1, 2025, and search results were imported into Rayyan software [17], a tool designed to facilitate the systematic screening of literature and duplicate removal. A detailed description of the database-specific search queries is provided in Supplementary Table 2.
2.3 Search Strategy
The search strategy was carefully designed to identify publications at the intersection of DP, DL, and medical applications. Specifically, the search strings combined three primary keyword groups: (1) differential privacy techniques (e.g., “differential privacy,” “differentially private,” “DP-SGD”), (2) deep learning and related machine learning methodologies (e.g., “deep learning,” “machine learning,” “neural network,” “artificial intelligence”), and (3) healthcare and clinical applications (e.g., “medical,” “clinical,” “healthcare,” “biomedical,” “EHR,” “imaging,” “radiology,” “text,” “speech”).
Each database query was adapted to database-specific indexing conventions to maximize search sensitivity. The comprehensive queries and exact search strings used, along with the number of articles retrieved from each database, are summarized in Supplementary Table 2. Additionally, manual searches were conducted by screening reference lists from key articles and reviewing recent preprints and related studies to capture any relevant works missed by the database searches.
2.4 Inclusion and Exclusion Criteria
The literature selection involved multiple screening steps: (1) Initial screening included automated duplicate removal and manual title and abstract screening in Rayyan. Articles were excluded if they lacked relevance to DP or healthcare applications. (2) Full-text reviews were conducted on the remaining articles to ensure alignment with inclusion criteria, specifically the empirical application of DP with DL techniques in medical contexts. Articles with purely theoretical content, lacking empirical evaluation, or applied to non-healthcare domains were excluded. The selection process ensured proportional representation across different medical application domains, including medical imaging, electronic health records, predictive analytics, and personalized medicine.
2.5 Synthesis of Results
The included studies were categorized using a structured two-step process. First, we grouped papers by their primary privacy mechanism, with a focus on applications of DP-SGD across clinical tasks and modalities, as summarized in Table 1. Design-related aspects—such as architectural choices, normalization strategies, and pretraining—are outlined in Table 2, while performance trends under varying privacy budgets are compiled in Table 3. Fairness evaluations, including subgroup analyses and reported metrics, are captured in Table 4. Studies employing alternative privacy mechanisms beyond DP-SGD are presented in Table 5, and empirical evaluations of attack resilience and defense strategies are detailed in Table 6.
3 Included studies
A total of 4,242 records were retrieved through systematic searches across four databases: PubMed, IEEE Xplore, ACM Digital Library, and Web of Science. After automated and manual removal of duplicates, 2,600 unique records remained. During the initial title and abstract screening phase, 2,529 articles were excluded based on the predefined inclusion and exclusion criteria, primarily due to lack of relevance to medical applications or absence of DP components.
The remaining 71 articles underwent full-text screening to assess their eligibility and methodological alignment with the review objectives. At this stage, additional exclusions were made based on insufficient empirical content, non-healthcare focus, or redundancy in covered topics, resulting in the exclusion of 20 more papers.
Ultimately, 51 studies met all criteria and were included from the database search. To supplement this, 23 additional relevant studies were identified through manual searching of reference lists and preprint repositories, ensuring coverage of the most recent and impactful contributions. In total, 74 papers were included in this scoping review. An overview of the entire screening and selection process is illustrated in Fig. 2. Additional descriptive characteristics of DP publications in medical AI are shown in Fig. 3.
Figure 2: Summary of study inclusion and descriptive statistics. (a) PRISMA flow diagram illustrating the study selection process for this scoping review. The diagram details the number of records identified through database and manual searches, duplicates removed, records screened by title and abstract, full-text articles assessed for eligibility, and the final number of studies included. (b) Publications per year on DP applications in medical DL, as of March 1, 2025. (c) Geographic distribution of included studies based on author affiliations; circle size reflects the number of publications with at least one affiliated author per country. Country outlines are based on Natural Earth.
Figure 3: Overview of study characteristics across included papers. Distribution of papers grouped by (a) modality or data type, (b) downstream prediction task, and (c) utilized neural network architectures. CNN: convolutional neural network; ECG: electrocardiogram; EEG: electroencephalography; EHR: electronic health record; GAN: generative adversarial network; GNN: graph neural network; MLP: multilayer perceptron; ViT: vision transformer.
4 Theoretical Background
We begin by establishing the core concepts and notation that form the basis of our discussion on DP in DL, followed by a detailed exposition of relevant privacy mechanisms and their integration into DL workflows.
Throughout this paper, we use uppercase bold letters (e.g., $\mathbf{X}$) for matrices, lowercase bold letters (e.g., $\mathbf{w}$) for vectors, uppercase letters (e.g., $D$) for datasets, and lowercase letters (e.g., $x$) for scalar variables or individual data samples.
4.1 Differential Privacy Definitions
DP was introduced by Dwork et al. [18] as a mathematically rigorous framework to limit the influence of any single individual’s data on the output of a computation. This protection is achieved by injecting controlled noise into a computation’s output, thereby masking the presence or absence of individual records.
Consider two neighboring datasets $D$ and $D^{\prime}$, differing by a single entry. A randomized algorithm $M$ is said to satisfy $(\varepsilon,\delta)$-DP if for all possible outputs $O$,
$\Pr[M(D)\in O]\leq e^{\varepsilon}\cdot\Pr[M(D^{\prime})\in O]+\delta$ (1)
Here, $\varepsilon$ controls the worst-case multiplicative difference in output probabilities between $D$ and $D^{\prime}$, while $\delta$ allows for a small additive probability of failure. When $\delta=0$, we recover the original, stricter definition of $\varepsilon$-DP.
In practice, $\delta$ is often set inversely proportional to dataset size ($\delta\sim 1/n$), ensuring that the likelihood of privacy failure remains negligible. Together, the $(\varepsilon,\delta)$ pair quantifies the privacy guarantee, with smaller values implying stronger protection.
Rényi differential privacy (RDP) [19] is a generalization of DP that leverages Rényi divergence [20] to provide tighter composition guarantees. The Rényi divergence of order $\alpha>1$ between the output distributions of $M$ on $D$ and $D^{\prime}$ is defined as:
$D_{\alpha}(M(D)\|M(D^{\prime}))=\frac{1}{\alpha-1}\log\mathbb{E}_{o\sim M(D^{\prime})}\left[\left(\frac{P_{M(D)}(o)}{P_{M(D^{\prime})}(o)}\right)^{\alpha}\right]$ (2)
A mechanism satisfies $(\alpha,\varepsilon)$-RDP if this divergence is bounded by $\varepsilon$ for all adjacent $D$ and $D^{\prime}$:
$D_{\alpha}\big{(}M(D)\,\|\,M(D^{\prime})\big{)}\leq\varepsilon$ (3)
Smaller values of $\alpha$ make the analysis more sensitive to small distributional differences, while larger values emphasize worst-case behavior. RDP can be converted to $(\varepsilon,\delta)$-DP using analytical bounds, and is widely used for accounting in iterative training.
The amount of noise required to achieve DP depends on the sensitivity of the function $f$ being computed, which quantifies how much $f$ can change when a single input is modified.
For vector-valued outputs, the $\ell_{2}$-sensitivity of $f$ is:
$\Delta_{2}f=\max_{\text{neighboring }D,D^{\prime}}\|f(D)-f(D^{\prime})\|_{2}$ (4)
This sensitivity bounds the required noise scale in mechanisms such as the Gaussian and Laplace mechanisms.
For real-valued outputs, the Gaussian mechanism adds noise $\eta\sim\mathcal{N}(0,\sigma^{2})$ to $f(D)$:
[…]
To achieve $(\varepsilon,\delta)$-DP, the standard deviation $\sigma$ must satisfy:
$\sigma\geq\frac{\Delta_{2}f\cdot\sqrt{2\ln(1.25/\delta)}}{\epsilon}$ (6)
For $\epsilon$-DP, the Laplace mechanism adds noise from $\mathrm{Lap}(0,b)$, where:
$b=\frac{\Delta_{1}f}{\epsilon}$ (7)
Here, $\Delta_{1}f$ is the $\ell_{1}$-sensitivity of $f$.
When the output is non-numeric (e.g., a selected category), the exponential mechanism selects an output $r$ from a range $\mathcal{R}$ using a utility function $u(D,r)$:
$\Pr[M(D)=r]\propto\exp\left(\frac{\epsilon\cdot u(D,r)}{2\Delta u}\right)$ (8)
with sensitivity:
$\Delta u=\max_{r\in\mathcal{R}}\max_{D,D^{\prime}}|u(D,r)-u(D^{\prime},r)|$ (9)
DP mechanisms can be applied at various stages of the learning pipeline, including the input [21, 22, 23], training process [11, 24], and model output [25, 26]. Among these, training-time DP has been shown to yield better accuracy and robustness [27].
4.2 Deep Learning and Optimization via SGD
DL refers to models that learn hierarchical representations from data via neural networks with multiple layers [28, 29]. Training these models involves minimizing an empirical risk objective:
$J(\boldsymbol{\theta})=\frac{1}{N}\sum_{i=1}^{N}L(\boldsymbol{\theta};x_{i},y_{i})$ (10)
where $L$ is the loss function and $\boldsymbol{\theta}$ the model parameters. This is typically optimized using stochastic gradient descent (SGD).
Variants of SGD differ by the batch size used in each update. In practice, mini-batch SGD is most common:
$\boldsymbol{\theta}_{t+1}=\boldsymbol{\theta}_{t}-\eta\cdot\frac{1}{m}\sum_{j=1}^{m}\nabla_{\boldsymbol{\theta}}L(\boldsymbol{\theta}_{t};x_{i_{j}},y_{i_{j}})$ (11)
where $\eta$ is the learning rate and $m$ is the mini-batch size. This formulation enables efficient training with good generalization.
4.3 Applying Privacy in Deep Neural Networks
Deep neural networks are vulnerable to privacy attacks targeting the input data, gradients, or model parameters [30]. Gradients are particularly sensitive, especially in federated settings where they are transmitted across devices [31]. Since gradients are computed directly from the input data, they can leak private information under adversarial scrutiny.
A common defense is to inject noise into gradients during training. DP-SGD [11] achieves this by modifying the gradient computation at each iteration. For each example $i$ in a mini-batch $\mathcal{B}$:
$\mathbf{g}_{i}$ $=\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};x_{i},y_{i})$ (12)
$\tilde{\mathbf{g}}_{i}$ $=\mathbf{g}_{i}\cdot\min\left(1,\frac{C}{\|\mathbf{g}_{i}\|_{2}}\right)$ (13)
$\bar{\mathbf{g}}$ $=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\tilde{\mathbf{g}}_{i}$ (14)
$\hat{\mathbf{g}}$ $=\bar{\mathbf{g}}+\mathcal{N}(0,\sigma^{2}C^{2}\mathbf{I})$ (15)
$\boldsymbol{\theta}$ $\leftarrow\boldsymbol{\theta}-\eta\hat{\mathbf{g}}$ (16)
Here, $C$ is the clipping norm and $\sigma$ is the noise multiplier. Fig. 4 illustrates this process. Clipping limits any individual’s contribution, while the added noise ensures DP guarantees.
Variants such as dynamic DP-SGD [32] adapt $C$ and $\sigma$ over time to improve utility during training. Other refinements target sparse gradients, which arise in applications like language models. A recent method [33] improves DP-SGD efficiency by focusing on non-zero gradient components, offering better performance in high-dimensional or sparse settings.
Figure 4: Overview of the DP-SGD training process. (a) In each iteration, a mini-batch of data is sampled. (b) Per-sample gradients are computed and clipped to a fixed norm to bound sensitivity. (c) Noise—typically Gaussian—is added to the aggregated clipped gradients to enforce differential privacy. (d) The model is updated using the noisy gradients. This process is repeated over multiple iterations, yielding a trained model with formal privacy guarantees.
4.4 Private Aggregation of Teacher Ensembles
Private aggregation of teacher ensembles (PATE), introduced by Papernot et al. [34], is a differentially private learning framework designed to protect sensitive training data by aggregating predictions from multiple independently trained models, referred to as ”teachers.” Each teacher model $T_{k}$ is trained on a disjoint subset $D_{k}$ of a sensitive dataset:
$D=\bigcup_{k=1}^{K}D_{k},\quad D_{i}\cap D_{j}=\emptyset\text{ for }i\neq j$ (17)
Given an input $x$, each teacher $T_{k}$ produces a class prediction $T_{k}(x)\in\{1,\dots,C\}$, and votes are tallied across the ensemble:
$v_{c}(x)=\sum_{k=1}^{K}\mathbb{I}[T_{k}(x)=c]$ (18)
To preserve privacy, Laplace noise is added to the vote counts:
$\tilde{v}_{c}(x)=v_{c}(x)+\text{Lap}(1/\lambda)$ (19)
Here, $\lambda$ controls the privacy level, with smaller values inducing stronger noise and stronger privacy.
A student model $S$ is then trained using public unlabeled data $D_{\text{public}}$, where noisy labels are obtained by querying the teacher ensemble. The final label assigned to each sample is:
$\hat{y}(x)=\arg\max_{c}\tilde{v}_{c}(x)$ (20)
The student is trained to minimize a supervised loss using the noisy labels:
$\mathcal{L}_{\text{student}}=\frac{1}{N}\sum_{i=1}^{N}\ell(S(x_{i}),\hat{y}(x_{i}))$ (21)
The total privacy loss is tracked using a privacy accountant, such as the moments accountant [11], ensuring that the cumulative privacy budget $\epsilon$ is not exceeded.
Compared to DP-SGD, which perturbs gradients within a single model during training, PATE decouples private learning from student training by using noise-perturbed knowledge transfer. PATE generally performs well in semi-supervised settings or when class imbalance can be mitigated via teacher ensembling. For example, a comparative study [35] showed that PATE-based GAN training often yields better utility-privacy tradeoffs than DP-SGD in cases where sensitive training data can be partitioned effectively.
4.5 Alternative Methods for Privacy Preservation
Beyond DP-SGD and PATE, several additional strategies have been developed to support privacy-preserving DL. One widely adopted framework is federated learning (FL), which enables collaborative model training across decentralized institutions or devices without requiring raw data to leave local storage [31]. FL can be implemented in centralized configurations—where local updates are sent to a coordinating server—or in decentralized topologies, where clients communicate directly with peers based on a predefined graph structure [36]. FL is increasingly used in clinical AI [4, 37, 38, 39], and is often combined with DP by adding noise to model updates. In such settings, privacy can be ensured either through centralized mechanisms—where noise is added after server aggregation—or local mechanisms, where clients inject noise prior to communication [40, 41].
Another major line of research involves secure multiparty computation, originally introduced by Yao [42]. It enables multiple parties to jointly compute functions over private inputs without revealing those inputs to each other. In DL, secure multiparty computation has been used to train or evaluate models securely across institutional boundaries [43]. Some recent works combine secure multiparty computation with DP [44, 45], adding noise either at each training iteration or once at the end of training. These hybrid designs differ in trade-offs: for instance, [44] focuses on communication efficiency through incremental protection, while [45] emphasizes accuracy preservation by applying noise post hoc.
Several methods aim to protect privacy by modifying gradients during training. These include clipping, pruning, masking, or injecting noise into gradients to obscure sensitive signals [46]. While some evidence suggests that such gradient masking can limit leakage without harming utility, other studies have shown that these defenses can be circumvented by more sophisticated adversaries [47].
An alternative line of defense involves perturbing internal feature representations rather than gradients. These feature-level defenses add noise or transformations at intermediate network layers, reducing the risk of inversion or memorization [48, 49, 50].
Synthetic data generation offers another privacy-preserving strategy. In this approach, a generative model (e.g., GAN or diffusion model) is trained on private data—sometimes under DP constraints—and is then used to produce synthetic datasets for downstream model training [51, 52]. This decouples the final model from direct exposure to sensitive records, reducing re-identification risk.
Classical data anonymization methods are still used in some clinical contexts. These include $k$-anonymity [53], $l$-diversity [54], and $t$-closeness [55], which aim to ensure that individual records are indistinguishable within equivalence classes. However, these approaches lack formal guarantees and are often vulnerable to auxiliary information attacks.
Another direction of research considers how architectural choices impact privacy. For example, empirical studies suggest that transformer-based models tend to leak more information than convolutional neural networks (CNNs), possibly due to higher capacity and attention mechanisms [56]. Components such as normalization layers, activation functions, and attention structures can influence a model’s vulnerability to privacy attacks.
Finally, cryptographic tools such as homomorphic encryption and secure aggregation are increasingly applied to protect data during training and communication [57]. In FL, for instance, gradients can be encrypted locally before transmission, adding an additional layer of protection against malicious servers or eavesdroppers.
These alternative mechanisms illustrate the diversity of tools available for protecting patient data in clinical machine learning. While our review focuses primarily on DP-SGD and PATE due to their theoretical foundations and empirical uptake, the broader ecosystem of privacy-preserving techniques—including encryption, secure computation, gradient manipulation, and architectural design—offers complementary strategies that may be better suited for specific tasks or deployment environments.
5 Applications of Differential Privacy in Clinical Deep Learning
5.1 Overview of Application Domains
DP has been applied to a wide range of clinical use cases, spanning diverse data modalities, deployment settings, and downstream tasks. Table 1 provides a summary of studies that applied DP via DP-SGD, the most widely adopted method in this domain. These studies encompass both centralized and FL settings, and cover modalities such as medical imaging (e.g., X-ray, CT, MRI), tabular data (e.g., EHR), biosignals (e.g., EEG, ECG), and multimodal inputs.
Study Modalities Application Downstream task Mech./account.
Adnan et al. [58] Histopathology FL Classification (lung cancer) GM/RDP
Al Aziz et al. [59] EHR Centralized EHR generation GM/MA
Chilukoti et al. [60] Histopathology Centralized Gastric cancer classification GM/RDP
Chin-Cheong et al. [61] EHR Centralized EHR generation GM/MA
Daum et al. [62] Cardiac cine MRI (4D) Centralized 3D MRI generation GM/RDP
DP-LPR [63] Genomic (DNA-SNPs) Centralized Gene interaction detection LM/NA
DP-SSLoRA [64] Chest X-ray Centralized Classification GM/RDP
DPFedSAM-Meas [65] Chest X-ray FL Classification GM, LM/RDP
Fan et al.[66] Tabular data Centralized Regression (survival analysis) GM/RDP
FedGAN [67] Chest X-ray FL Classification GM/NA
Guo et al. [68] Dermoscopic photography Centralized Classification GM/NA
Hatamizadeh et al. [69] Chest X-ray, fundoscopy FL Classification GM/NA
Jones et al. [70] Time series Centralized Classification GM/RDP
Kaess et al. [71] CT Centralized Detecting CT contrast agents GM/NA
Khanna et al. [72] Genomic data FL Classification GM/RDP
Lal et al. [73] Tabular data Centralized Classification of fetal status GM/MA
Mehmood et al. [74] MRI FL Classification (brain tumor, AD) GM/NA
Mueller et al. [75] Image-derived graph Centralized Whole-graph classification GM/RDP
Odeyomi et al. [76] Chest X-ray Centralized Pneumonia classification GM/RDP
P3SGD [77] Histopathology Centralized Classification GM/MA
Pan et al. [78] Histopathology FL Segmentation GM/NA
PP-LDG [79] CT Centralized Segmentation (liver tumors) GM/MA
Shiri et al. [80] Chest CT FL COVID-19 prognosis GM/NA
Tayebi Arasteh et al. [13] X-ray Centralized Classification GM/RDP
Tayebi Arasteh et al. [5] X-ray, abdominal CT Centralized Classification GM/RDP
Tayebi Arasteh et al. [81] Pathological speech Centralized Neurological disorder analysis GM/RDP
Torfi et al. [82] EEG, EHR, ECG, biopsy Centralized Data generation GM/RDP
Wang et al. [83] Tabular data FL Classification GM/MA
Ziller et al. [84] X-ray, CT, fundoscopy Centralized Classification, segmentation GM/GDP
Table 1: Summary of studies applying DP-SGD in medical deep learning. Each entry includes the data modality, learning setting (centralized or FL), downstream task, and the DP mechanism and accounting method used. Only studies using DP-SGD are included. AD: Alzheimer’s disease; ECG: electrocardiography; EEG: electroencephalography; EHR: electronic health record; FL: federated learning; GM: Gaussian mechanism; LM: Laplace mechanism; MA: moments accountant; NA: information not available; RDP: Rényi differential privacy.
5.1.1 Modalities and Clinical Contexts
The reviewed studies apply DP-SGD across a broad range of medical data types, with imaging emerging as the most common modality. Chest X-rays dominate classification tasks [13, 5, 64, 67, 84, 65, 79, 71], while MRI—particularly brain and cardiac cine MRI—is used in both classification and generative modeling applications [74, 62]. Histopathology appears in several studies focusing on cancer classification and segmentation [58, 78, 77, 60].
Beyond imaging, DP-SGD has been applied to structured data such as EHRs and tabular datasets [66, 73, 61, 59], genomic sequences [63, 72], biosignals including EEG and ECG [82], and pathological speech [81]. Time series data is also explored in classification contexts [70], and some generative frameworks integrate multimodal data such as EEG, biopsy, and EHR [82].
Classification remains the predominant downstream task [13, 5, 64, 67, 84, 65, 71, 74, 60, 66, 73, 72, 68, 70], followed by generative modeling [62, 61, 59, 82]. Despite growing interest, applications involving longitudinal data or multimodal fusion remain underexplored.
5.1.2 Deployment Settings: Centralized vs Federated Learning
DP-SGD is employed in both centralized [13, 5, 64, 63, 66, 84, 75, 76, 73, 68, 79, 77, 62, 81, 61, 59, 60, 71, 82, 70] and federated [80, 83, 67, 72, 69, 58, 74, 65, 78] training pipelines. In centralized settings, models are trained on institutional data and privatized using DP-SGD, protecting against post-deployment leakage. These include imaging [13, 5, 64, 84], structured/tabular data [66, 73], and EHRs [61, 59].
FL-based applications involve multi-institutional collaboration, where local clients update models on-site and share only privatized gradients. These setups target imaging [80, 67, 69], genomics [72], and histopathology [58, 78]. Despite their promise, FL studies often omit key DP parameters, complicating reproducibility.
While FL offers inherent privacy through data decentralization, centralized DP-SGD remains vital for models intended for public sharing or external evaluation. The versatility of DP-SGD across both paradigms underscores its relevance to privacy-preserving clinical AI.
5.1.3 Mechanisms and Privacy Accounting for DP-SGD
The Gaussian mechanism (GM), paired with Rényi differential privacy (RDP), is the dominant configuration in DP-SGD implementations, offering favorable composition properties and practical support through libraries like Opacus and TensorFlow Privacy [85]. Most reviewed works adopt GM with RDP for per-sample gradient perturbation.
Nonetheless, reporting practices are inconsistent. Some studies provide complete details—including $(\epsilon,\delta)$ values, clipping norms, and noise multipliers [13, 5, 66, 75, 81]—while others omit the accounting mechanism or final budget [80, 67, 69]. A few exceptions use the Laplace mechanism (LM) [63] or combine GM and LM [65]. The moments accountant (MA) is used in some centralized and FL settings [83, 73].
Privacy budgets range from conservative ($\epsilon\approx 1$) to more relaxed ($\epsilon\approx 10$). However, few studies systematically explore multiple $\epsilon$ levels, limiting our understanding of utility degradation under stronger privacy guarantees.
Overall, while GM and RDP dominate, standardization in reporting mechanisms, parameters, and accounting tools remains a critical need for enabling rigorous comparisons and reproducibility across studies.
5.2 Design Choices for Privacy-Utility Tradeoffs
Introducing DP into DL workflows inevitably introduces tradeoffs between model utility and privacy guarantees. The severity of this tradeoff is highly sensitive to several key design decisions, including architecture choice, normalization strategy, and data augmentation. Table 2 summarizes these technical configurations across reviewed studies.
Study Modalities Architectures Normalization Data aug.
Adnan et al. [58] Lung cancer detection MEM-MIL NA NA
Beaulieu-Jones et al. [70] Time series AC-GAN [86] NA NA
Chin-Cheong et al. [61] EHR Wasserstein GAN MSN NA
Chilukoti et al. [60] Histopathology NF-Net [87] NA NA
Daum et al. [62] Cardiac cine MRI (4D) Latent diffusion models[88] NA NA
Fan et al. [66] Tabular data MLP NA NA
Guo, Yu et al. [68] Dermoscopic images NA GN NA
Kaess et al. [71] CT ResNet-9 Scale norm NA
Khanna et al. [72] Genomic data MLP NA NA
Mehmood et al. [74] MRI InceptionV3 [89] NA Affine
Mueller et al. [75] Image-derived graph Graph neural networks Instance norm NA
Nguyen et al.[67] Chest X-ray GAN NA NA
Riess et al. [39] MRI NA GN NA
Tayebi Arasteh et al. [13] Chest X-ray ResNet-9 GN None
Tayebi Arasteh et al. [5] Chest X-ray, abdominal CT ResNet-9 [90] GN None
Tayebi Arasteh et al. [81] Pathological speech ResNet-18 GN None
Yan et al. [64] Chest X-ray ResNet-18 None NA
Zhang et al. [63] Genomic (DNA-SNPs) Custom CNN NA NA
Ziller et al. [84] Chest X-ray, Liver CT VGG-11, Custom U-Net None Affine
Table 2: Model design strategies for managing privacy-utility tradeoffs in DP-SGD applications. The table lists architectures, data normalization techniques, and data augmentation practices used in studies applying DP-SGD to medical deep learning. GN: group normalization; MEM-MIL: multi-expert MIL; MLP: multilayer perceptron; MSN: mode-specific normalization; NA: information not available.
5.2.1 DL Architectures for DP-SGD
Architecture selection significantly influences a model’s resilience to the noise and clipping operations imposed by DP-SGD. Compact CNNs such as ResNet-9 and ResNet-18 were the most widely adopted across classification tasks in chest X-ray, CT, and speech data [13, 5, 64, 81, 71]. Other lightweight backbones like VGG-11 and custom U-Net variants were applied in medical imaging tasks [84], while NF-Net was used in histopathology [60], and InceptionV3 in brain MRI [74].
Generative models used GAN architectures, including Wasserstein GANs for EHR [61], AC-GANs for time series [70], and standard GANs in FL imaging tasks [67]. More recent approaches explored latent diffusion models for cardiac MRI synthesis [62]. MLPs remain prevalent for tabular and genomic data [66, 72], while specialized networks such as graph neural networks (GNNs) [75] and MEM-MIL [58] were applied in niche contexts like graph classification and histopathology.
Notably, vision transformers (ViTs) [91] and other large-scale models were absent from this set, echoing findings in separate comparisons that such architectures are more sensitive to DP noise and tend to underperform under strong privacy constraints.
5.2.2 Normalization Techniques for DP-SGD
Batch normalization (BN), which depends on mini-batch statistics, is generally avoided in DP-SGD because it interferes with per-sample gradient computation. Among studies reporting normalization explicitly, group normalization (GN) was most common, used in both classification and speech tasks [13, 5, 81, 68, 39]. GN avoids batch dependencies and offers more stable training under DP noise.
Other strategies include instance normalization for graph neural networks [75], mode-specific normalization (MSN) for generative EHR models [61], and scale normalization in CT classification [71]. Several studies—particularly those using GANs, MLPs, or CNNs in tabular or genomic settings—did not specify any normalization approach [70, 63, 66, 67, 72, 58]. Others reported disabling normalization entirely [64, 84], likely to avoid incompatibility with gradient clipping.
Overall, there is no universal best practice, though GN appears robust across diverse tasks. Systematic benchmarking of normalization techniques under DP constraints remains a key direction for future research.
5.2.3 Data Augmentation Practices for DP-SGD
Data augmentation, a staple of conventional DL, is rarely applied under DP-SGD. Among reviewed studies, only two used basic affine transformations—one for brain MRI classification [74] and one for multi-modal segmentation [84]. Most others either explicitly avoided augmentation [13, 5, 81] or did not report it.
This reluctance stems from the risk of increasing gradient variance, which exacerbates training instability under noisy DP updates. Augmentations can alter input distributions in ways that interfere with the gradient signal, especially in small medical datasets with high class imbalance [5].
The lack of systematic evaluation limits conclusions about which augmentations, if any, are compatible with DP-SGD. There is a clear need for controlled experiments testing standard and domain-specific augmentation strategies under DP conditions.
5.2.4 Pretraining and Transfer Learning for DP-SGD
Pretraining is a consistent and effective strategy for mitigating performance loss under DP-SGD. By starting from pretrained weights, models avoid the fragile convergence dynamics associated with noisy training from scratch.
Tayebi Arasteh et al. [5] demonstrated that DP-SGD models trained from scratch failed to learn meaningful features, whereas pretrained ResNet-9 models retained high accuracy. In a related study, pretrained models on chest X-ray and CT were shown to maintain near non-private performance under tight $\epsilon$ budgets [5]. Pretraining on general-domain speech corpora also improved DP training on pathological speech classification [81], despite cross-domain mismatch. Similarly, ImageNet-initialized models improved performance in brain MRI classification under DP [74].
Pretraining is not only helpful but often necessary for training deeper or more expressive models under DP-SGD. Explicit reporting and standardization of pretraining strategies is vital for reproducibility and benchmarking.
5.3 Privacy-Utility Tradeoff Results
A core challenge in differentially private DL is navigating the trade-off between model utility and privacy. As summarized in Table 3, studies vary widely in how they report performance under different privacy budgets, with only a subset evaluating multiple $\epsilon$ levels. This inconsistency limits comparative understanding across clinical modalities and tasks.
Study Modality Non-private Very low budget Low budget
Adnan et al. [58] Histopathology Accuracy = $82\%$ $\epsilon=2.9$/Accuracy = $81\%$ $\epsilon=10$/Accuracy = $78\%$
[…]
Alsenani et al. [93] Sensor data Accuracy = $30\%$ NA $\epsilon=10$/Accuracy = $27\%$
Beaulieu-Jones et al. [70] Time series Spearman = $0.96$ $\epsilon=3$/Spearman = $0.91$ NA
Chin-Cheong et al. [61] EHR AUC = $80.0\%$ $\epsilon=1$/AUC = $66.1\%$ $\epsilon=10$/AUC = $67.8\%$
Daum et al. [62] Cardiac cine MRI FID = $15.4$ $\epsilon=1$/FID = $29.8$ $\epsilon=10$/FID = $26.8$
Fan et al. [66] Tabular data IBS = ~$17$ $\epsilon=1$/IBS = ~$21$ $\epsilon=8$/IBS = ~$18$
Fu et al. [94] EEG & sensor Accuracy = $99.3\%$ $\epsilon=1$/Accuracy = $98.7\%$ NA
Gong et al. [95] Time series Accuracy = $88.5\%$ $\epsilon=2$/Accuracy = $75.3\%$ $\epsilon=10$/Accuracy = $78.2\%$
Guo, Yu et al. [68] Dermoscopy Accuracy = $69.1\%$ $\epsilon=3$/Accuracy = ~$43\%$ $\epsilon=10$/Accuracy = $61.2\%$
Kaess et al. [71] CT Accuracy = $99.5\%$ NA $\epsilon=8$ /Accuracy = $97.4\%$
Mehmood et al. [74] MRI Accuracy = $88\%$ NA $\epsilon=5.6$/Accuracy = $78\%$
Ming Y et al. [96] Tabular EMR DWS =$0.0041$ $\epsilon=1$/DWS = $0.05$ $\epsilon=10$/DWS = $0.0061$
Mueller et al. [75] Image-derived graph AUC = $98.5\%$ $\epsilon=1$/AUC = $95.1\%$ $\epsilon=10$/AUC = $97.2\%$
Nguyen et al. [67] Chest X-ray Accuracy = $99\%$ $\epsilon=0.3$/Accuracy = ~$90\%$ NA
Odeyomi et al. [97] Chest X-ray Accuracy = $92\%$ $\epsilon=2$/Accuracy = $88\%$ $\epsilon=8$/Accuracy = $90\%$
Riess et al. [39] MRI Accuracy = $90.1\%$ $\epsilon=1$/Accuracy = $81.6\%$ $\epsilon=10$/Accuracy = $90\%$
Sun et al. [98] Tabular data DOP = ~$0.75$ $\epsilon=1$/DOP = ~$1.2$ $\epsilon=10$/DOP = ~$1.2$
Tang et al. [99] Tabular data F1-score = $19.6\%$ $\epsilon=1$/F1-score = $19.79\%$ NA
Tayebi Arasteh et al. [13] Chest X-ray AUC$=68-90\%$ $\epsilon\approx 1$/AUC$=67-89\%$ $\epsilon\approx 10$/AUC$=67-89\%$
Tayebi Arasteh et al. [5] Chest X-ray AUC$=89.7\%$ $\epsilon=0.5$/AUC$=84.0\%$ $\epsilon=7.9$/AUC$=87.4\%$
Tayebi Arasteh et al. [5] CT AUC$=99.7\%$ $\epsilon=0.5$/AUC$=92\%$ $\epsilon=8.0$/AUC$=99.3\%$
Tayebi Arasteh et al. [81] Pathological speech Accuracy = $99.1\%$ $\epsilon=0.9$/Accuracy = $88.3\%$ $\epsilon=7.5$/Accuracy = $95.3\%$
Wang et al. [100] Sensor data AUC = $59\%$ $\epsilon=1$/AUC = $56\%$ NA
Yan et al. [64] Chest X-ray AUC = $99\%$ $\epsilon=2$/AUC = $96.5\%$ $\epsilon=10$/AUC = $98\%$
Zhang et al. [63] DNA-SNPs Accuracy = $96\%$ $\epsilon=1$/Accuracy = ~$87\%$ NA
Zhang et al. [101] Chest X-ray Accuracy = $92.9\%$ $\epsilon=1$/Accuracy = $84.7\%$ $\epsilon=10$/Accuracy = $91.5\%$
Ziller et al. [84] Chest X-ray AUC = $96\%$ $\epsilon=0.5$/AUC = $84\%$ NA
Ziller et al. [84] Liver CT AUC = $96\%$ $\epsilon=0.5$/Dice= $94\%$ NA
Table 3: Reported privacy-utility tradeoffs across varying privacy budgets. This table summarizes model performance in both non-private settings and under different levels of protection, focusing on very strong protection (very low: $\epsilon\approx 1$) and strong protection (low: $\epsilon\approx 5$–10). Note that some studies applied DP to more than one modality or task; in such cases, each modality-task pair is listed as a separate row. Metrics vary by task and include accuracy, area under the receiver operating characteristic curve (AUC), F1-score, Spearman correlation, integrated Brier score (IBS), distance of optimal calibration (DOP), distance between Wasserstein barycenters (DWS), Dice score, and Fréchet inception distance (FID), as reported by each study. NA: Information not available.
5.3.1 Performance at Extremely Strong Privacy Levels
At very low privacy budgets ($\epsilon\approx 1$), the magnitude of injected noise often leads to significant performance degradation. For example, Tayebi Arasteh et al. [5] observed an AUC drop from 89.7% to 84.0% on chest X-rays and from 99.7% to 92.0% on abdominal CT when applying $\epsilon=0.5$. In pathological speech classification, accuracy fell from 99.1% to 88.3% at $\epsilon=0.9$ [81], and in dermoscopy, accuracy declined from 69.1% to  43% at $\epsilon=3$ [68]. Similar trends were reported for generative tasks, such as cardiac MRI synthesis, where Fréchet inception distance (FID) worsened substantially under $\epsilon=1$ [62].
However, there are exceptions. In a multi-institutional study on domain transfer, Tayebi Arasteh et al. [13] showed that external validation performance remained nearly unchanged across $\epsilon\approx 1$–10. This suggests that DP noise may act as a regularizer, benefiting generalization even as in-domain performance drops.
5.3.2 Utility Under Strong Privacy Levels: A Modality-Specific Perspective
Most reviewed studies operate in the strong privacy regime ($\epsilon\approx 5$–10), where performance generally remains close to non-private baselines. For instance, chest X-ray AUC dropped modestly from 89.7% to 87.4% at $\epsilon=7.9$, and CT classification remained above 99% even at $\epsilon=8$ [5]. In MRI [74] and genomic data [72], performance losses at $\epsilon=10$ were similarly modest.
Nonetheless, more fragile domains show greater vulnerability to DP noise. Studies in histopathology [58], dermoscopy [68], and speech [81] report larger performance drops, reflecting higher input variance, smaller datasets, and more complex learning tasks. Tabular and sensor-based applications—e.g., in mental health or activity monitoring—also showed significant utility loss at lower $\epsilon$ values [92, 100].
These trends emphasize the importance of modality-specific tuning: modalities like chest imaging are more robust under DP-SGD, while others require additional regularization, pretraining, or architectural adaptation to maintain clinical viability.
While most studies report $\epsilon$, the value of $\delta$—which bounds the probability of privacy failure—is often missing or inconsistently stated. Best practice recommends $\delta\leq 1/n$, where $n$ is the dataset size. Many studies follow this, setting $\delta$ between $10^{-5}$ and $10^{-6}$, but others omit it or use heuristic values without justification.
Reporting gaps extend beyond privacy parameters. Performance under multiple $\epsilon$ levels is rarely benchmarked, and core implementation details—such as clipping norm, noise multiplier, accounting method (e.g., RDP or moments accountant), and number of training epochs—are frequently missing. This limits reproducibility and cross-study comparison.
This issue is particularly acute in FL settings, where DP is often applied to distributed model updates. Although FL introduces additional privacy through decentralization, few studies evaluate how DP noise affects utility across varying budgets in such settings, hindering generalizability.
To address these gaps, we recommend several reporting standards for future work. (1) Researchers should report both non-private and private model performance to provide clear context for evaluating the impact of privacy mechanisms. (2) The values of key privacy parameters, including $\epsilon$, $\delta$, and the accounting method used (e.g., RDP or moments accountant), should be explicitly stated to ensure interpretability and comparability. (3) Performance metrics should be reported across multiple $\epsilon$ levels to characterize the trade-off between privacy and utility. (4) Finally, studies should document core model and training details, including architecture type, dataset size, pretraining strategies, and other relevant hyperparameters. Adopting these practices will improve transparency, foster reproducibility, and support more rigorous benchmarking of privacy-preserving models for clinical deployment.
6 Fairness Considerations in DP-Based Medical DL
6.1 Privacy-Fairness Interactions
While DP is designed to protect individual data, its application in DL can unintentionally exacerbate performance disparities across demographic subgroups [12, 5, 102]. This is particularly concerning in healthcare, where systemic biases already exist, and privacy mechanisms that disproportionately impact underrepresented groups can further entrench inequities. Table 4 summarizes reviewed studies that assess the fairness implications of DP in medical DL, the subgroup attributes evaluated, and the metrics used.
Study Modalities Use case Fairness groups
DP privacy-fairness
Chin-Cheong et al. [61] EHR EHR generation Age, sex
Kaess et al. [71] Chest CT Classification Race, age, sex
Tayebi Arasteh et al. [13] Chest X-ray Classification Age, sex
Tayebi Arasteh et al. [5] Chest X-ray, abdominal CT Classification Age, sex
Tayebi Arasteh et al. [81] Pathological speech Neurological disorder classification Age, sex
Fairness metrics
Average odds difference [103] Any Mean gap in TPR and FPR across groups Any
Disparate impact ratio[104, 105] Any Ratio of positive prediction rates Any
Equal accuracy [103] Any Difference in overall accuracy across groups Any
Equal opportunity [103] Any Difference in TPR across groups Any
Equalized odds difference [103] Any Difference in both TPR and FPR Any
FDR parity [106] Any Difference in FDR across groups Any
FOR parity [106] Any Difference in FOR across groups Any
Group-wise calibration [107] Any Of predicted probabilities Any
Overdiagnosis rate [108] Imaging, tabular Difference in false positive rates across groups Age, sex
Predictive equality [103] Any Difference in FPR across groups Any
PtD [109, 5] Any Positive prediction rate gap across groups Any
Subgroup AUC [5] Any, especially imaging AUC reported per group to assess disparities Age, sex, race
Treatment equality [110] Any Ratio of FN to FP across groups Any
Underdiagnosis rate [108] Imaging, tabular Difference in false negative rates across groups Age, sex
Table 4: Summary of privacy-fairness studies and fairness evaluation metrics. The upper section lists studies that explicitly investigate fairness impacts of applying DP in medical deep learning, including use cases and demographic groups analyzed. The lower section summarizes fairness metrics reported across studies, covering various dimensions of group-level equity. ”Any” in the modalities or fairness groups column indicates that the metric is applicable across any modality or protected attribute (e.g., age, sex, race). AUC: area under the receiver operating characteristic curve; EHR: electronic health record; FDR: false discovery rate; FOR: false omission rate; FPR: false positive rate; PtD: statistical parity difference; TPR: true positive rate.
Only a handful of studies explicitly examine fairness under DP constraints [61, 71, 13, 5, 81]. These focus primarily on group-level disparities in classification tasks across attributes such as age, sex, and race. For instance, Tayebi Arasteh et al. evaluated fairness across sex and age in chest X-ray [5] and speech classification [81]; Kaess et al. [71] examined racial disparities in chest CT; and Chin-Cheong et al. [61] analyzed synthetic EHR outputs.
Findings suggest that the impact of DP is highly context-dependent. In some cases, privacy noise interacts with data imbalance to disproportionately degrade performance for certain groups—especially minorities with fewer examples. For example, DP constraints in speech classification disproportionately harmed age groups [81].
Despite these insights, most studies report fairness at only a single privacy level. Only a few works—such as [13, 5, 81]—systematically analyze fairness across multiple $\epsilon$ values, limiting our understanding of how fairness evolves as privacy constraints tighten.
Moreover, fairness is rarely a design objective. While certain configurations (e.g., ResNet backbones or group normalization) were found to reduce subgroup variance [5, 81], these choices were primarily utility-driven. Fairness-aware techniques like constraint-based learning, adversarial debiasing, or subgroup reweighting are almost entirely absent.
Broader gaps remain. Most evaluations consider only binary attributes, overlooking intersectional identities (e.g., older female patients), and focus almost exclusively on classification, neglecting regression and survival tasks. Additionally, the lack of demographically diverse benchmarks hinders fairness auditing under DP constraints. Progress will require treating fairness as a first-class design goal in DP pipelines, and systematically reporting subgroup performance across varying privacy levels.
6.2 Fairness Metrics and Their Application
Fairness auditing is essential when applying DP to medical DL, as noise can distort performance differently across groups [111]. Reviewed studies primarily assess group-level metrics based on protected attributes such as age, sex, or race. Table 4 summarizes the main metrics used.
Most fairness evaluations focus on group-level comparisons, typically assessing disparities across sex, age, or race. The most frequently reported metric is statistical parity difference (PtD) [109, 5, 13, 81, 112, 113], which measures the difference in the proportion of positive predictions between two subgroups:
$\text{PtD}=\Pr(\hat{Y}=1\mid A=a)-\Pr(\hat{Y}=1\mid A=b)$ (22)
Here, $\hat{Y}$ is the model prediction, $A$ is a protected attribute (e.g., sex), and $a$, $b$ are two group values (e.g., male and female). A related metric is the disparate impact ratio (DIR) [104, 105], which compares the rates of favorable outcomes as a ratio:
$\text{DIR}=\frac{\Pr(\hat{Y}=1\mid A=a)}{\Pr(\hat{Y}=1\mid A=b)}$ (23)
Another widely used metric is the equal opportunity difference [103, 81], which focuses on differences in true positive rates (TPRs) between groups:
$\Delta\text{EqOp}=\Pr(\hat{Y}=1\mid Y=1,A=a)-\Pr(\hat{Y}=1\mid Y=1,A=b)$ (24)
In this case, $Y$ is the ground truth label, with $Y=1$ indicating a positive case. Equalized odds [103] extends this concept by considering both true positive and false positive rates:
$\Delta\text{EOD}=$ $\left|\Pr(\hat{Y}=1\mid Y=1,A=a)-\Pr(\hat{Y}=1\mid Y=1,A=b)\right|$
$+\left|\Pr(\hat{Y}=1\mid Y=0,A=a)-\Pr(\hat{Y}=1\mid Y=0,A=b)\right|$ (25)
Predictive equality [103], by contrast, focuses only on disparities in false positive rates:
$\Delta\text{FPR}=\Pr(\hat{Y}=1\mid Y=0,A=a)-\Pr(\hat{Y}=1\mid Y=0,A=b)$ (26)
Treatment equality [110] compares the balance between false negatives (FN) and false positives (FP) across groups:
Treatment $\text{Treatment Equality}=\frac{\text{FN}_{a}/\text{FP}_{a}}{\text{FN}_{b}/\text{FP}_{b}}$ (27)
In addition to general-purpose fairness metrics, several studies employed domain-specific measures such as underdiagnosis rate [108] (differences in false negative rates across groups), overdiagnosis rate (false positives), and subgroup AUC [5, 81, 13]. While these do not require formal equations, they are particularly important in clinical settings, where different types of errors carry different consequences for patient care.
While these metrics are useful, they must be interpreted in context. For example, parity in outcomes across age groups may be inappropriate when prevalence rates differ biologically. Furthermore, DP-induced noise may distort probability calibration, reducing the reliability of threshold-based or probabilistic fairness measures.
To ensure robustness, future studies should adopt a standardized approach to fairness auditing under DP: (1) report multiple fairness metrics to capture different dimensions of disparity; (2) justify the choice of protected attributes based on clinical relevance; (3) disaggregate results by subgroup across multiple $\epsilon$ values; and (4) apply calibration or post hoc adjustments when using probability-based metrics. The development of benchmark datasets with detailed demographic annotations is also critical to advance fair, privacy-preserving AI in medicine.
7 Alternative Privacy Mechanisms Beyond DP-SGD
7.1 Federated and Mechanism-Agnostic Privacy Approaches
While DP-SGD remains the most widely used method for enforcing DP in medical DL, it is not always ideal for real-world clinical scenarios. Centralized training can be infeasible when data is decentralized, and DP-SGD often struggles with convergence or model stability under strict privacy budgets. As a result, a growing body of work explores privacy mechanisms beyond DP-SGD—including local differential privacy (LDP), exponential and Laplace mechanisms, secure aggregation protocols, and hybrid federated designs. Table 5 summarizes these approaches across various clinical tasks and data modalities.
Study Modalities Downstream task Mech./account.
Al Aziz et al. [92] Tabular data Classification LM, EM/LC
Almadhoun et al. [114] Genomic data Genomic privacy protection LM/NA
Alsenani et al. [93] Sensor data Detect patient activity patterns LM, GM, EM/LC
Ay et al. [115] Brain MRI Brain tumor classification NA/NA
Chen et al. [116] DNA-Seq DNA motif discovery RR/NA
Díaz et al. [117] Brain MRI AD classification GM/NA
Faisal et al. [118] Chest X-ray Pneumonia classification GM/RDP
Fan et al. [94] EEG, sensor data Freezing of gait detection LM/LC
Fang et al. [119] Tabular data synthetic data generation GM/MA
Fedeli et al. [120] Tabular data Survival prediction LM/NA
Gong et al. [95] Time series Decision support in ICUs GM/MA
Gwon et al. [96] EHR Secure data sharing RR, LM/NA
Hatamizadeh et al. [99] Tabular data Heart failure prediction NA/NA
He et al. [121] RNA-Seq Disease vs. non-disease classification GM/LC
He et al. [122] Genomic data Synthetic genomic data generation LM/NA
Honkela et al. [123] RNA-Seq Drug sensitivity prediction LM/LC
Kim et al. [124] Sensor data Health monitoring LM/LC
Kong et al. [125] Histopathology Prostate cancer diagnosis GM/NA
Li et al. [126] Pathological speech Classification (AD detection) LM, GM/NA
Li et al. [127] EHR Classification (diagnosis) LM/NA
Ming Y et al. [128] Histopathology Classification, survival prediction GM/NA
Movahedi et al. [129] Tabular data Classification LM, EM/NA
Niinimäki et al. [130] Genomic data Classification, regression LM/LC
Sun et al. [98] Tabular data Data generation GM/RDP
Tchouka et al. [131] Unstructured clinical text Multi-label ICD-10 code classification LM, EM/NA
Tramèr et al. [132] Genomic data Identifying significant SNPs LM, EM/NA
Veeraragavan et al. [133] EHR Survival Analysis LM/LC
Wang et al. [100] Sensor data Stress detection LM/NA
Wei et al. [134] Tabular data Heart disease prediction GM/LC
Wu et al. [135] Tabular data Secure data sharing RR/NA
Yuan et al. [97] Chest X-ray Pneumonia classification GM/GDP
Zhang et al. [101] Chest X-ray Pneumonia diagnosis GM/NA
Zhang et al. [136] EHR Release of medical datasets LM/LC
Table 5: Summary of studies applying alternative DP mechanisms beyond DP-SGD. Each entry includes the data modality, downstream task, and the DP mechanism and accounting method used. These include applications of local DP, encryption-based aggregation, federated learning (FL) variants, and generative modeling frameworks across a range of clinical modalities and tasks. AC: advanced composition; AD: Alzheimer’s disease; EHR: electronic health record; EM: exponential mechanism; GDP: Gaussian differential privacy; GM: Gaussian mechanism; PD: Parkinson’s disease; NA: information not available; LC: linear composition; LM: Laplace mechanism; MA: moments accountant; PLD: privacy loss distribution; RDP: Rényi DP accountant; RR: randomized response.
Several studies extended DP to FL settings, where collaborative training is performed without centralizing patient data. Alsenani et al. [93] introduced the FAItH framework, which combines local and centralized DP with secure multiparty computation for sensor data clustering. Gong et al. [95] used Gaussian noise with moment accounting in a federated intensive care unit monitoring system. Yuan et al. [97] employed Gaussian DP in a federated chest X-ray classification pipeline. Riess et al. [39] adapted federated training for MRI classification using complex-valued neural networks.
Some studies (e.g., [100, 115]) demonstrated FL-based medical models under DP constraints but did not report full accounting details. This lack of standardization highlights ongoing challenges in reproducibility and transparency.
Outside of FL, many studies used alternative mechanisms in centralized or generative contexts. Laplace mechanisms were used in tasks such as pathological speech classification [126], cancer risk modeling [129], and drug sensitivity prediction [123]. Exponential mechanisms were applied in EHR and clinical text classification tasks [131, 92], often in combination with Laplace noise or local perturbation.
Some studies, such as Wu et al. [135], adopted one-shot local DP strategies where perturbation occurs per sample, requiring no composition accounting. Similarly, Gwon et al. [96] used randomized response mechanisms for synthetic EHR generation, emphasizing privacy at the point of data release rather than during model training.
Generative modeling also featured prominently. He et al. [121] and Zhang et al. [136] synthesized RNA-Seq and EHR data under DP. Notably, Sun et al. [98] applied noise to the loss gradients—rather than individual per-sample gradients—when training a generator, a setup inspired by the DP-GAN framework. This contrasts with DP-SGD-based GANs (e.g., [82]) that apply DP to per-sample gradients in the discriminator.
Other applications included DP for data sharing [136], motif discovery in DNA sequences [116], and survival analysis using exponential and Laplace mechanisms [120, 133].
Despite their diversity, alternative mechanisms often suffer from inconsistent reporting. Many studies omit essential parameters, such as the privacy budget $(\epsilon)$, the delta value, or the type of accountant used. Mechanisms like randomized response or one-step perturbation—while avoiding the need for composition—still require careful tuning to balance utility and privacy. Encryption-based designs introduce additional complexity through communication overhead and synchronization requirements.
Nevertheless, these approaches highlight a shift toward more flexible, mechanism-aware privacy strategies. By leveraging differentially private techniques suited to specific tasks and data types, these methods offer scalable alternatives to DP-SGD—particularly in federated, generative, or low-resource settings. Moving forward, clearer guidelines for reporting, accounting, and mechanism selection will be essential for advancing fair, privacy-preserving AI in healthcare.
7.2 Defenses Against Attacks
As privacy-preserving DL becomes increasingly integrated into clinical pipelines, adversarial attacks against these models have grown in sophistication. DP provides a formal framework for limiting such risks, but the real-world effectiveness of DP depends on implementation details, perturbation granularity, and the threat model assumed. Table 6 summarizes empirical evaluations of DP-enabled defenses across various attack types in healthcare applications.
Study Perturbation Attack type Utilized DP
Almadhoun et al. [114] GP Attribute inference attack, MIA DP-SGD
Almadhoun et al. [114] GP MIA DP-SGD
Bingzhe et al. [77] GP Model-inversion attack DP-SGD
Díaz et al. [117] Weight perturbation Client inference attack Beyond DP-SGD
DPFL [80] GP MIA DP-SGD
Fan et al. [66] GP MIA DP-SGD
Gwon et al. [96] Input perturbation MIA Beyond DP-SGD
Hatamizadeh et al. [69] GP GIA (BN-aware inversion) DP-SGD
Hyunwook et al. [137] Model-based perturbation Re-identification risk assessment Beyond DP-SGD
Kaissis et al. [4] GP Model-inversion attack DP-SGD
Liangrui et al. [78] GP GIA DP-SGD
Riess et al. [39] GP GIA DP-SGD
Shunrong et al. [65] GP MIA DP-SGD
Sun et al. [98] GP Identity disclosure attack DP-SGD
Sun et al. [98] GP Attribute disclosure attack DP-SGD
Tramèr et al. [132] Output perturbation MIA (theoretical) Beyond DP-SGD
Veeraragavan et al. [133] GP MIA DP-SGD
You et al. [93] Output perturbation MIA Beyond DP-SGD
Ziller et al. [111] GP Reconstruction attack DP-SGD
Table 6: Summary of attack evaluations on DP models in healthcare. Each row indicates the perturbation method used (e.g., gradient or input perturbation), the type of attack evaluated (e.g., membership inference, gradient inversion), and whether the study used DP-SGD or alternative DP methods. GIA: gradient inversion attack; GP: gradient perturbation; IP: input perturbation; MIA: membership inference attack.
Membership inference attacks (MIAs) attempt to infer whether a specific sample was part of the model’s training dataset—a particularly serious risk in healthcare, where inclusion in a disease cohort may itself reveal sensitive clinical information. Multiple studies applied DP-SGD to protect against MIAs in classification tasks using imaging, sensor, and tabular data [80, 66, 65, 133]. These defenses typically introduced Gaussian noise into gradient updates, reducing attack success rates even at moderate $\epsilon$ values.
However, the strength of protection varied across settings. Fan et al. [66] noted that class imbalance and shallow architectures increased MIA vulnerability under DP, emphasizing the need to consider model and data characteristics when tuning privacy parameters. Almadhoun et al. [114] confirmed that even with DP-SGD, certain MIAs could still succeed under weak noise or limited clipping, highlighting the importance of robust DP configurations.
Beyond MIAs, several studies evaluated inversion-based threats, where attackers attempt to reconstruct sensitive inputs—such as medical images or EHR records—by analyzing shared gradients. These attacks are especially relevant in federated learning, where gradient exchange is common. For example, Hatamizadeh et al. [69] and Liangrui et al. [78] demonstrated that DP-SGD defenses could substantially degrade inversion quality but not eliminate leakage entirely. Ziller et al. [111] observed similar results in reconstruction attacks on multi-institutional imaging models.
Alternative perturbation schemes, such as weight noise [117] or complex-valued gradients [39], were also explored as mitigation strategies. Still, these methods remain under-tested and lack standardized evaluation criteria, especially under stronger adversary assumptions (e.g., white-box gradient access).
LDP, which perturbs data before model access, was investigated in a few studies (e.g., [93, 96, 132]). While theoretically strong, LDP was shown to be insufficient in practice against sophisticated attacks, including attribute inference and partial MIAs. These findings underscore that DP defenses must often extend beyond input perturbation, especially in distributed learning settings.
A recurring recommendation across studies is the use of hybrid defenses—combining DP with cryptographic methods like secure aggregation or multiparty computation [93, 135]. These layered strategies offer stronger guarantees but introduce practical constraints, including communication overhead and increased system complexity.
Despite growing interest in attack resilience, several gaps remain. First, the majority of studies focus on a narrow set of attacks—most commonly MIAs—while overlooking others like training-time poisoning or client-level inference in federated setups. Second, many reports lack standardized threat models or omit key experimental details (e.g., batch size, gradient access scope), limiting reproducibility. Third, defenses are rarely evaluated under varying $\epsilon$ levels, making it difficult to assess tradeoffs between privacy strength and robustness.
DP offers a foundational layer of protection against common inference attacks in medical DL. However, achieving effective defense requires not only tuning $\epsilon$, but also understanding how model architecture, data distribution, and perturbation method interact with adversarial goals. Future work should prioritize comprehensive, reproducible evaluations that benchmark attack resilience under clinically realistic settings and privacy constraints.
8 Discussion
This scoping review synthesizes recent advances in applying DP to medical DL, with a particular focus on subgroup fairness, attack resilience, utility tradeoffs, and alternative privacy mechanisms. Across 74 studies, we observed growing maturity in technical implementations, but uneven adoption of standardized evaluation protocols and fairness-aware design practices.
Despite the theoretical strengths of DP [10], practical implementations face well-documented tradeoffs. Performance consistently degrades under stricter privacy budgets, particularly at $\epsilon\approx 1$, where noise injection frequently leads to convergence failure or sharp accuracy declines [5, 68, 81]. However, several studies demonstrated that with strong pretraining [5, 74], stable architectures [13, 5, 71], and careful normalization [81], models can maintain clinically acceptable utility even under strong privacy guarantees (ϵitalic-$\epsilon\approx$10). Imaging tasks, particularly chest X-ray and CT classification, showed resilience to DP noise [13, 5], while domains such as histopathology [58], dermoscopy [68], and pathological speech [81] proved more vulnerable, largely due to smaller datasets or higher input variability.
Fairness remains underexplored relative to utility. A handful of studies—most notably those by Tayebi Arasteh et al. [13, 5, 81]—conducted thorough evaluations of subgroup disparities under varying privacy levels, revealing that DP can exacerbate demographic gaps unless explicitly mitigated. However, fairness is rarely treated as a core training objective. Intersectional effects are largely absent, and fairness in regression or survival tasks remains rarely assessed [66, 133]. Fairness-aware optimization methods such as reweighting or constraint-based training are virtually unexplored in the current literature. Metrics like subgroup AUC, underdiagnosis rate, and statistical parity difference have been used to track disparities [13, 71, 61], but no consensus exists on which are most clinically meaningful or how to interpret them under noisy predictions.
Alternative privacy mechanisms beyond DP-SGD—including Laplace and exponential noise [126, 92, 120], federated aggregation [93, 95, 97], and local perturbation [96, 135]—offer increased flexibility and often better alignment with deployment constraints. Generative models, such as GANs and diffusion models, were applied to EHR and medical image synthesis under DP constraints [121, 119, 98]. However, many studies omit essential parameters like $(\epsilon,\delta)$ or the privacy accountant used, complicating reproducibility and comparison.
We also reviewed the threat landscape for privacy-preserving models. Empirical evaluations show that DP can significantly mitigate attacks such as membership inference and gradient inversion [80, 66, 69, 111], but protections are highly context-dependent. Fan et al. [66] emphasized that class imbalance and shallow architectures increased vulnerability. Almadhoun et al. [114] demonstrated residual attack success under weak noise, even with DP-SGD. Some studies also evaluated reconstruction attacks in FL [78, 138], identity and attribute disclosure risks [98], and defenses using weight perturbation [117] or complex-valued networks [39]. Importantly, LDP alone was shown to be insufficient under adversarial conditions [96, 132].
This review has several limitations. While our search covered major venues and was supplemented by manual inclusion of key preprints, the rapid pace of publication means that recent work may have been missed. In addition, while we systematically included all eligible studies applying DP to medical DL, some clinical areas—such as genomics [63, 130] or survival modeling [66, 133]—are still underrepresented in the literature itself. Finally, our review focused primarily on technical methods and performance outcomes; perspectives from clinicians, patients, and data custodians were outside the scope but remain essential for translating privacy-preserving models into practice.
Several future directions emerge. First, fairness evaluations should become standard practice in DP studies, ideally spanning multiple $\epsilon$ levels and well-justified subgroup definitions. Second, benchmark datasets with demographic annotations and support for intersectional analysis would facilitate robust, reproducible fairness evaluations. Third, new methods that integrate privacy and fairness directly into training—such as group-aware DP optimizers or adaptive clipping—could help ensure that protections do not exacerbate health inequities. Fourth, clearer documentation of DP parameters, accountant types, and performance tradeoffs will improve transparency and support rigorous benchmarking.
Differential privacy offers a robust framework for protecting individual data in clinical AI models. However, privacy alone does not guarantee equitable or reliable deployment. A shift toward integrated design—where privacy, fairness, and utility are optimized together—is essential for developing trustworthy, real-world-ready medical AI. This review provides a foundation for that goal, clarifying what is already known and identifying what remains to be addressed.
Additional information
Author contributions: The formal analysis was conducted by MM and STA. The original draft was written by MM and STA. The literature search was performed by STA, and the screening was performed by MM, MV, ML, and STA. MR, DT, and STA provided clinical expertise. MM, MV, ML, MR, DT, AM, and STA provided technical expertise. STA defined the study. Figures were designed by MM. All authors read the manuscript, contributed to the editing, and agreed to the submission of this paper.
Ethical approval: No human or animal subjects are involved in this study.
List of Abbreviations
  • AD - Alzheimer’s disease
  • AUC - Area under the receiver operating characteristic curve
  • BN - Batch normalization
  • DL - Deep learning
  • DP - Differential privacy
  • DP-SGD - Differentially-private stochastic gradient descent
  • ECG - Electrocardiogram
  • EEG - Electroencephalography
  • EHR - Electronic health record
  • EM - Exponential mechanism
  • FDR - False discovery rate
  • FID - Fréchet inception distance
  • FL - Federated learning
  • FOR - False omission rate
  • FPR - False positive rate
  • GAN - Generative adversarial network
  • GIA - Gradient inversion attack
  • GM - Gaussian mechanism
  • GN - Group normalization
  • LM - Laplace mechanism
  • LDP - Local differential privacy
  • MA - Moments accountant
  • MIA - Membership inference attack
  • MSN - Mode-specific normalization
  • PATE - Private aggregation of teacher ensembles
  • PD - Parkinson’s disease
  • PRISMA - Preferred reporting items for systematic reviews and meta-analyses
  • PRISMA-ScR - PRISMA extension for scoping reviews
  • PtD - Statistical parity difference
  • RDP - Rényi differential privacy
  • SGD - Stochastic gradient descent
  • TPR - True positive rate
  • ViT - Vision transformer
Appendix A Supplementary information
Table S.1: PRISMA-ScR checklist [14]
Section Item PRISMA-ScR checklist item Status
Title
Title 1 Identify the report as a scoping review.
Abstract
Structured summary 2 Provide a structured summary that includes (as applicable): background, objectives, eligibility criteria, sources of evidence, charting methods, results, and conclusions that relate to the review questions and objectives.
Introduction
Rationale 3 Describe the rationale for the review in the context of what is already known. Explain why the review questions/objectives lend themselves to a scoping review approach.
Objectives 4 Provide an explicit statement of the questions and objectives being addressed with reference to their key elements (e.g., population or participants, concepts, and context) or other relevant key elements used to conceptualize the review questions and/or objectives.
Methods
Protocol and registration 5 Indicate whether a review protocol exists; state if and where it can be accessed (e.g., a Web address); and if available, provide registration information, including the registration number.
Eligibility criteria 6 Specify characteristics of the sources of evidence used as eligibility criteria (e.g., years considered, language, and publication status), and provide a rationale.
Information sources 7 Describe all information sources in the search (e.g., databases with dates of coverage and contact with authors to identify additional sources), as well as the date the most recent search was executed.
Search 8 Present the full electronic search strategy for at least 1 database, including any limits used, such that it could be repeated.
Selection of sources of evidence 9 State the process for selecting sources of evidence (i.e., screening and eligibility) included in the scoping review.
Data charting process 10 Describe the methods of charting data from the included sources of evidence (e.g., calibrated forms or forms that have been tested by the team before their use, and whether data charting was done independently or in duplicate) and any processes for obtaining and confirming data from investigators.
Data items 11 List and define all variables for which data were sought and any assumptions and simplifications made.
Critical appraisal of individual sources of evidence 12 If done, provide a rationale for conducting a critical appraisal of included sources of evidence; describe the methods used and how this information was used in any data synthesis (if appropriate). NA
Synthesis of results 13 Describe the methods of handling and summarizing the data that were charted.
Results
Selection of sources of evidence 14 Give numbers of sources of evidence screened, assessed for eligibility, and included in the review, with reasons for exclusions at each stage, ideally using a flow diagram.
Characteristics of sources of evidence 15 For each source of evidence, present characteristics for which data were charted and provide the citations
Critical appraisal within sources of evidence 16 If done, present data on critical appraisal of included sources of evidence (see item 12). NA
Results of individual sources of evidence 17 For each included source of evidence, present the relevant data that were charted that relate to the review questions and objectives.
Synthesis of results 18 Summarize and/or present the charting results as they relate to the review questions and objectives
Discussion
Summary of evidence 19 Summarize the main results (including an overview of concepts, themes, and types of evidence available), link to the review questions and objectives, and consider the relevance to key groups.
Limitations 20 Discuss the limitations of the scoping review process
Conclusions 21 Provide a general interpretation of the results w.r.t the review questions and objectives, as well as potential implications and/or next steps.
Funding
Funding 19 Describe sources of funding for the included sources of evidence, as well as sources of funding for the scoping review. Describe the role of the funders of the scoping review.
Database Results Search string
PubMed 63 (”differential privacy”[TIAB] OR ”differentially private”[TIAB] OR ”DP-SGD”[TIAB]) AND (”deep learning”[TIAB] OR ”machine learning”[TIAB] OR ”neural network”[TIAB]) AND (”medicine”[TIAB] OR ”medical”[TIAB] OR ”clinical”[TIAB] OR ”biomedical”[TIAB] OR ”health”[TIAB] OR ”EHR”[TIAB] OR ”imaging”[TIAB] OR ”radiology”[TIAB] OR ”text”[TIAB] OR ”speech”[TIAB]) NOT (review[PT])
84 (”differential privacy” OR ”differentially private” OR ”DP-SGD”) AND (”deep learning” OR ”machine learning” OR ”neural network” OR ”artificial intelligence”) AND (”medical” OR ”medicine” OR ”clinical” OR ”biomedical” OR ”healthcare” OR ”EHR” OR ”imaging” OR ”radiology” OR ”MRI” OR ”CT” OR ”ultrasound” OR ”speech” OR ”text” OR ”video”) NOT (review[PT])
417 ( ”differential privacy”[TIAB] OR ”differentially private”[TIAB] OR ”DP-SGD”[TIAB] OR ”privacy preserving”[TIAB] OR ”privacy-preserving”[TIAB] OR ”securing”[TIAB] OR ”safeguarding”[TIAB] ) AND ( ”deep learning”[TIAB] OR ”machine learning”[TIAB] OR ”artificial intelligence”[TIAB] OR ”neural network”[TIAB] ) AND ( ”medical”[TIAB] OR ”medicine”[TIAB] OR ”clinical”[TIAB] OR ”healthcare”[TIAB] OR ”biomedical”[TIAB] OR ”EHR”[TIAB] OR ”electronic health record”[TIAB] OR ”imaging”[TIAB] OR ”radiology”[TIAB] OR ”text”[TIAB] OR ”speech”[TIAB] ) NOT (review[PT])
IEEE Xplore 676 (”differential privacy” OR ”differentially private” OR ”DP-SGD”) AND (”deep learning” OR ”machine learning” OR ”neural network”) AND (”medicine” OR ”medical” OR ”clinical” OR ”healthcare” OR ”biomedical” OR ”EHR” OR ”imaging” OR ”radiology” OR ”text” OR ”speech”)
Web of Science 494 ”differential privacy” AND (”deep learning” OR ”machine learning”) AND (medical OR medicine OR clinical OR healthcare OR biomedical OR imaging OR EHR OR radiology OR text OR speech)
571 (”differential privacy” OR ”differentially private” OR ”DP-SGD”) AND (”deep learning” OR ”machine learning” OR ”artificial intelligence” OR ”neural network”) AND (medical OR medicine OR clinical OR healthcare OR biomedical OR imaging OR EHR OR radiology OR text OR speech)
ACM Digital Library 999 ”differential privacy” AND (”deep learning”) AND (medical OR EHR OR imaging)
938 ”differential privacy” AND (”artificial intelligence”) AND (medical OR clinical OR healthcare OR EHR OR imaging OR radiology OR biomedical)
Table S.2: Search results from different databases